# Stabilizing Non-Abelian Topological Order against Heralded Noise via Local Lindbladian Dynamics

### v1

@@ Abstract
An important open question for the current generation of highly controllable quantum devices is understanding which phases can be realized as stable steady-states under local quantum dynamics. In this work, we show how robust steady-state phases with both Abelian and non-Abelian mixed-state topological order can be stabilized against generic “heralded” noise using active dynamics that incorporate measurement and feedback, modeled as a fully local Lindblad master equation. These topologically ordered steady states are two-way connected to pure topologically ordered ground states using local quantum channels, and preserve quantum information for a time that is exponentially large in the system size. Specifically, we present explicit constructions of families of local Lindbladians for both Abelian ($\mathbb{Z}_{2}$) and non-Abelian ($D_{4}$) topological order whose steady-states host mixed-state topological order when the noise is below a threshold strength. As the noise strength is increased, these models exhibit first-order transitions to intermediate mixed state phases where they encode robust classical memories, followed by (first-order) transitions to a trivial steady state at high noise rates. When the noise is imperfectly heralded, steady-state order disappears but our active dynamics significantly enhances the lifetime of the encoded logical information. To carry out the numerical simulations for the non-Abelian $D_{4}$ case, we introduce a generalized stabilizer tableau formalism that permits efficient simulation of the non-Abelian Lindbladian dynamics.
@@ I Introduction
Topological phases of matter are characterized by a number of intriguing properties, including multiple locally indistinguishable ground states with subtle entanglement structure that enables them to store quantum information. However, while these phases are stable to local perturbations of the Hamiltonian that do not close the many-body gap [1, 2, 3], in $d\leq 3$, topological order does not survive at any nonzero temperature [4, 5, 6, 7, 8, 9, 10, 11]. [^1] This is fundamentally due to the presence of deconfined point-like excitations which, when thermally excited, destroy the long-range entanglement in the ground-space (even if the excitations have restricted mobility, as in fracton models [12, 13, 14]). However, certain models in $d\geq 4$ (e.g., the 4d Toric code [15, 16]) lack any point-like excitations and are thermally stable phases [17, 8].
The fact that even a very cold environment is incompatible with low-dimensional TO raises the question of whether such orders can be stabilized at all in noisy open quantum systems. One approach to achieving this in two spatial dimensions (2d) is quantum error correction [15, 18, 19, 20, 21, 22, 23, 24, 25], which, when performed at regular intervals, can protect information stored in the ground space of a topologically ordered Hamiltonian (which we henceforth refer to as a TO state) for arbitrarily long times. More specifically, in a TO state that is subjected to sufficiently weak noise for a sufficiently short time, a round of error-syndrome measurements, followed by a (typically nonlocal) recovery operation, is sufficient to fully recover the original state. Conversely, if the noise is too strong or acts for too long time, information about the initial state is irretrievably lost in an information-theoretic phase transition. The existence of such mixed-state phase transitions has led to a surge of interest in investigating topologically non-trivial mixed states that result from local noise channels [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49] and in defining notions of mixed-state phases of matter [50, 51, 52, 53, 54, 55, 56].
@@ II Background and Summary of Main Results
Figure 1: Steady-state phase diagrams for the two models studied, with active dynamics removing $X$ and $Z$ flags at rates $\gamma_{x}$ and $\gamma_{z}$, respectively. The heralded depolarizing noise rate is set to $\eta=1$. The color indicates the combined density $\frac{1}{2}(n^{f}_{X}+n^{f}_{Z})$ of flags evaluated at a late time ($t_{f}=100$ for the Toric code and $t_{f}=20$ for the $D_{4}$ model) at which the densities (for parameters away from the phase boundaries) have saturated to time-independent values (see Appendix A). Each data point is obtained by averaging over $100$ independent Monte Carlo realizations. (a) Phase diagram for the Toric code on a $48\times 48$ honeycomb lattice, showing an active phase where $(n^{f}_{X}+n^{f}_{Z})/2$ remains small in the steady state, a partially $X$ ($Z$) active phase where $n^{f}_{X}$ ($n^{f}_{Z}$) remains small in the steady state, while $n^{f}_{Z}\ (n^{f}_{X})=1$, and an absorbing state where $n^{f}_{X}=n^{f}_{Z}=1$. The active phase has $\mathbb{Z}_{2}$ steady-state TO, while the partially active phase has an intrinsically mixed TO. (b) Phase diagram for the non-Abelian $D_{4}$ model defined on three interpenetrating $48\times 48$ sized honeycomb lattices, exhibiting an active phase with steady-state $D_{4}$ TO, a partially $X$-active phase, and a topologically trivial absorbing phase. The active phases protect the initially encoded quantum information up to exponentially long times, while the partially active phases only protect classical information for long times.
@@
The protocol we consider can be written as a Lindblad master equation $\partial_{t}\rho=\mathcal{L}(\rho)$ with two types of jump operators, corresponding respectively to the heralded noise and the recovery procedure. We have specified the recovery procedure as a “measurement-and-feedback” process; however, any jump operator can be written in this form through a polar decomposition. We are interested in the steady-state manifolds of the Lindblad superoperator $\mathcal{L}$ as a function of the noise and recovery rates. Steady-states are eigenvectors of $\mathcal{L}$ with zero eigenvalue. In a stable TO phase, we expect $\mathcal{L}$ to have a nontrivial manifold of states with eigenvalues that are exponentially close in modulus to 00 (in linear system size).
@@
An important caveat is in order: although we characterize phases in terms of their steady-state order (or lack thereof), in any finite-sized system there is always a non-vanishing probability that the system ends up in an absorbing configuration (i.e., with all erasure flags occupied). Once the dynamics lands in this configuration, the system cannot escape it; hence, in any finite-sized system, strictly speaking the $t\to\infty$ steady-state is always the absorbing state. However, the probability of reaching this fixed-point state can be exponentially small in the system size $L$ [77], such that the time-scale $\tau$ at which the absorbing state is reached increases exponentially with $L$. When we refer to a steady-state in this regime, we will thus always mean the steady-state of the system up to these exponentially long time-scales (which are beyond the range of our numerical simulations for all system sizes considered). Correspondingly, by a non-trivally ordered steady-state phase we mean one that does not reach the absorbing state for times that are exponentially long in the system size. On the other hand, there exist regimes in parameter space where the absorbing state is reached on a time-scale (for large enough $L$) that is independent of system size. We refer to such regimes, for which the absorbing state is the only possible choice of steady-state, as being in the absorbing phase.
@@
To illustrate our protocol in a simple setting, we first consider the ground-space of the 2d Toric code subject to both local correction dynamics and unbiased heralded errors; the resulting steady-state phase diagram is shown in Fig. 1(a). Our local correction protocol succeeds in stabilizing a topologically non-trivial mixed steady-state (the active phase in the Figure) for noise below a certain threshold. Specifically, we show that this mixed-state (a) preserves logical information by applying the minimum weight perfect matching (MWPM) decoder, and (b) is connected to the original pure state by a local quantum channel whose depth scales at most logarithmically with the number of qubits. Thus, the active phase of our dynamics belongs to a mixed-state phase with $\mathbb{Z}_{2}$ TO. We also find that a first-order phase transition separates this active phase from neighbouring dynamical phases, of which two (the $X$-active and $Z$-active phases in the Figure) preserve only classical information [^3] These phases are characterized by the $\mathbb{Z}_{2}^{(0)}$ anyon theory and, in the parlance of Ref. [40], constitute an intrinsically mixed TO..
@@
We then turn to our main focus, which is to show that local correction dynamics succeeds in stabilizing a mixed steady-state phase with $D_{4}$ TO for perfectly heralded noise (below a certain threshold). $D_{4}$ TO belongs to the general class of TOs which harbour point-like non-Abelian anyons: whereas Abelian anyons only acquire a universal phase when braided around each other, the internal degeneracy of non-Abelian anyons can be used to construct a richer set of quantum gates, which can be exploited for topological quantum computation [73, 78, 79, 80]. While states with arbitrary non-Abelian TO are challenging to prepare, those derived from solvable groups [^4] Given a finite group $G$, its derived series is a set of normal subgroups $N_{j}$, derived inductively via $N_{0}=G$ and $N_{j}=[N_{j-1},N_{j-1}]$ ($j>0$). Here, $[N,N]$ denotes the commutator subgroup of $N$. The derived length of $G$ is defined as the smallest positive integer $\ell_{G}$ such that $N_{\ell_{G}}=\mathbb{Z}_{1}$. A solvable group is defined as one for which $\ell_{G}$ is finite. permit efficient preparation via adaptive quantum circuits [81, 82, 83, 84] and have been realized on NISQ devices [85, 86, 64, 87, 88]. In the specific case of $D_{4}$ TO, non-Abelian braiding was recently demonstrated on a trapped-ion processor [64].
@@
Finally, a technical contribution of this work is a generalization of the stabilizer tableau [92, 93] to the case of “quasi-stabilizers” (defined below), which permits efficient simulation of Lindblad dynamics for a certain class of non-Pauli stabilizers.
@@ III.1 Noise model for unbiased heralded errors
We first introduce the noise model that we will use throughout this work, which consists of unbiased, heralded errors. To model a system with heralded errors, in addition to the qubits themselves, we also introduce a classical degree of freedom, which we refer to as an erasure “flag”. These flags are represented by a pair of occupation numbers $n^{f}_{X_{j}}$ and $n^{f}_{Z_{j}}$ which are associated with each qubit $j$ and can be in either occupied ($n^{f}=1$) or unoccupied ($n^{f}_{X}=0$) state. Here, we will set all flags to be unoccupied $n^{f}_{X_{j}}=n^{f}_{Z_{j}}=0\,\forall\,j$ in the initial state.
@@
Here, $j$ labels the qubit index and $\mathsf{q_{x},q_{z}}\in\{0,1\}$. The operators $f^{+}_{X_{j}}(f^{-}_{X_{j}})$ raise $n^{f}_{X_{j}}$ from 00 to $1$ (lower $n^{f}_{X_{j}}$ from $1$ to 00). The relation $f^{+}_{X_{j}}f^{-}_{X_{j}}=n^{f}_{X_{j}}$ implies that the jump operator labeled by $\mathsf{q_{x}}=1(\mathsf{q_{x}}=0)$ acts on the site if an $X$ flag was already present (absent). The operators $f^{\pm}_{Z_{j}}$ act analogously on the $Z$ flags. Here

### v2

@@ Abstract
An important open question for the current generation of highly controllable quantum devices is understanding which phases can be realized as stable steady-states under local quantum dynamics. In this work, we show how robust steady-state phases with both Abelian and non-Abelian mixed-state topological order can be stabilized, in two spatial dimensions (2d), against generic “heralded” noise using active dynamics that incorporate measurement and feedback, modeled as a fully local Lindblad master equation. These topologically ordered steady states are two-way connected to pure topologically ordered ground states using local quantum channels, and preserve quantum information for a time that is exponentially large in the system size. Specifically, we present explicit constructions of families of local Lindbladians for both Abelian ($\mathbb{Z}_{2}$) and non-Abelian ($D_{4}$) topological order whose steady-states host mixed-state topological order when the noise is below a threshold strength. As the noise strength is increased, these models exhibit first-order transitions to intermediate mixed state phases where they encode robust classical memories, followed by (first-order) transitions to a trivial steady state at high noise rates. When the noise is imperfectly heralded, steady-state order disappears but our active dynamics significantly enhances the lifetime of the encoded logical information. To carry out the numerical simulations for the non-Abelian $D_{4}$ case, we introduce a generalized stabilizer tableau formalism that permits efficient simulation of the non-Abelian Lindbladian dynamics.
@@ I Introduction
Topological phases of matter are characterized by a number of intriguing properties, including multiple locally indistinguishable ground states with subtle entanglement structure that enables them to store quantum information. However, while these phases are stable to local perturbations of the Hamiltonian that do not close the many-body gap [1, 2, 3], in $d\leq 3$, topological order (TO) does not survive at any nonzero temperature [4, 5, 6, 7, 8, 9, 10, 11]. [^1] This is fundamentally due to the presence of deconfined point-like excitations which, when thermally excited, destroy the long-range entanglement in the ground-space (even if the excitations have restricted mobility, as in fracton models [12, 13, 14]). However, certain models in $d\geq 4$ (e.g., the 4d Toric code [15, 16]) lack any point-like excitations and are thermally stable phases [17, 8].
The fact that even a very cold environment is incompatible with low-dimensional TO raises the question of whether such orders can be stabilized at all in noisy open quantum systems. One approach to achieving this in two spatial dimensions (2d) is quantum error correction [15, 18, 19, 20, 21, 22, 23, 24, 25, 26], which, when performed at regular intervals, can protect information stored in the ground space of a topologically ordered Hamiltonian (which we henceforth refer to as a TO state) for arbitrarily long times. More specifically, in a TO state that is subjected to sufficiently weak noise for a sufficiently short time, a round of error-syndrome measurements, followed by a (typically nonlocal) recovery operation, is sufficient to fully recover the original state. Conversely, if the noise is too strong or acts for too long time, information about the initial state is irretrievably lost in an information-theoretic phase transition. The existence of such mixed-state phase transitions has led to a surge of interest in investigating topologically non-trivial mixed states that result from local noise channels [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51] and in defining notions of mixed-state phases of matter [52, 53, 54, 55, 56, 57, 58, 59].
@@ II Background and Summary of Main Results
Figure 1: Steady-state phase diagrams for (a) the Toric code on a $48\times 48$ honeycomb lattice and (b) the non-Abelian $D_{4}$ model defined on three interpenetrating $48\times 48$ sized honeycomb lattices. The active dynamics removes the $X$ and $Z$ flags at rates $\gamma_{x}$ and $\gamma_{z}$, respectively. The heralded depolarizing noise rate is set to $\eta=1$. The background color indicates the combined density $\frac{1}{2}(n^{f}_{X}+n^{f}_{Z})$ of flags evaluated at a late time ($t_{f}=100$ for the Toric code and $t_{f}=20$ for the $D_{4}$ model) at which the densities (for parameters away from the phase boundaries) have saturated to time-independent values (see Appendix A). The logical error rates are computed for a select set of parameters depicted using the diamond-shaped symbols. These are colored according to $1-\frac{1}{2}(F_{|\widetilde{0}\rangle_{L}}+F_{|\widetilde{+}\rangle_{L}})$, which quantifies the steady-state logical error rate averaged over two distinct initial logical states. Each data point is obtained by averaging over $100$ independent Monte Carlo realizations. For both models, the active phase of the flags ($n^{f}_{X},n^{f}_{Z}\approx 0$) leads to a topologically ordered steady-state phase, where both logical fidelities $F$ are close to $1$. This phase protects the initially encoded quantum information up to exponentially long times. In contrast, the partially active phases (only one type of flags has small density) can only protect classical information for long times, signaled by only one of the $F\approx 1$. The detailed time evolution for representative points in each of these phases, highlighted using cyan colored squares, is shown in Fig. 5 and $1$.
@@
The protocol we consider can be written as a Lindblad master equation $\partial_{t}\rho=\mathcal{L}(\rho)$ with two types of jump operators, corresponding respectively to the heralded noise and the recovery procedure. We have specified the recovery procedure as a “measurement-and-feedback” process; however, any jump operator can be written in this form through a polar decomposition. We are interested in the steady-state manifolds of the Lindblad superoperator $\mathcal{L}$ as a function of the noise and recovery rates. Steady-states are eigenvectors of $\mathcal{L}$ with zero eigenvalue. In a stable TO phase, we expect $\mathcal{L}$ to have a nontrivial manifold of states with eigenvalues that are exponentially close in modulus to 0 (in linear system size).
@@
An important caveat is in order: although we characterize phases in terms of their steady-state order (or lack thereof), in any finite-sized system, there is always a non-vanishing probability that the system ends up in an absorbing configuration (i.e., with all erasure flags occupied). Once the dynamics lands in this configuration, the system cannot escape it; hence, in any finite-sized system, strictly speaking the $t\to\infty$ steady-state is always the absorbing state. However, the probability of reaching this fixed-point state can be exponentially small in the system size $L$ [80], such that the time-scale $\tau$ at which the absorbing state is reached increases exponentially with $L$. When we refer to a steady-state in this regime, we will thus always mean the steady-state of the system up to these exponentially long time-scales (which are beyond the range of our numerical simulations for all system sizes considered). Correspondingly, by a non-trivially ordered steady-state phase we mean one that does not reach the absorbing state for times that are exponentially long in the system size. On the other hand, there exist regimes in parameter space where the absorbing state is reached on a time-scale (for large enough $L$) that is independent of system size. We refer to such regimes, for which the absorbing state is the only possible choice of steady-state, as being in the absorbing phase.
@@
The central concept of our protocol is that it leverages information about the locations of stabilizer defects, provided by the heralded noise channel and encoded in the classical flags, to effectively confine the stabilizer defects using fully local dynamics. As a result of the competition between our local correction protocol and the heralded noise, the classical flags undergo an absorbing state transition from an active phase to the absorbing phase. The flag densities ($n^{f}_{X}$ and $n^{f}_{Z}$) serve as an order parameter for such transitions [81, 80]. Moreover, we show that the small values of the flag densities in the active phase precisely correspond to topological order in the steady state, as directly evidenced by low logical error rates in this phase.
@@
To illustrate our protocol in a simple setting, we first consider the ground-space of the 2d Toric code subject to both local correction dynamics and unbiased heralded errors; the resulting steady-state phase diagram is shown in Fig. 1(a). Our local correction protocol succeeds in stabilizing a topologically non-trivial mixed steady-state (the active phase in Fig. 1) for noise below a certain threshold. Specifically, we show that this mixed-state (a) preserves logical information by applying the minimum weight perfect matching (MWPM) decoder, and (b) is connected to the original pure state by a local quantum channel whose depth scales at most logarithmically with the number of qubits. Thus, the active phase of our dynamics belongs to a mixed-state phase with $\mathbb{Z}_{2}$ TO. We also find that a first-order phase transition separates this active phase from neighboring dynamical phases, of which two (the $X$-active and $Z$-active phases in Fig. 1(a)) preserve only classical information [^3] These phases are characterized by the $\mathbb{Z}_{2}^{(0)}$ anyon theory and, in the parlance of Ref. [41], constitute an intrinsically mixed TO..
@@
We then turn to our main focus, which is to show that local correction dynamics succeeds in stabilizing a mixed steady-state phase with $D_{4}$ TO for perfectly heralded noise (below a certain threshold). $D_{4}$ TO belongs to the general class of TOs which harbor point-like non-Abelian anyons: whereas Abelian anyons only acquire a universal phase when braided around each other, the internal degeneracy of non-Abelian anyons can be used to construct a richer set of quantum gates, which can be exploited for topological quantum computation [76, 82, 83, 84]. While states with arbitrary non-Abelian TO are challenging to prepare, those derived from solvable groups [^4] Given a finite group $G$, its derived series is a set of normal subgroups $N_{j}$, derived inductively via $N_{0}=G$ and $N_{j}=[N_{j-1},N_{j-1}]$ ($j>0$). Here, $[N,N]$ denotes the commutator subgroup of $N$. The derived length of $G$ is defined as the smallest positive integer $\ell_{G}$ such that $N_{\ell_{G}}=\mathbb{Z}_{1}$. A solvable group is defined as one for which $\ell_{G}$ is finite. permit efficient preparation via adaptive quantum circuits [85, 86, 87, 88] and have been realized on NISQ devices [89, 90, 67, 91, 92]. In the specific case of $D_{4}$ TO, non-Abelian braiding was recently demonstrated on a trapped-ion processor [67]. Motivated in part by these experiments, we focus on TO in $d=2$ spatial dimensions in this paper. While we expect in principle that our protocol extends to $d=3$ spatial dimensions mutatis mutandis (since it is the point-like anyons that destabilize the phase, just as in $d=2$), we do not pursue this here since the numerical simulations are far more challenging.
@@
A technical contribution of this work is a generalization of the stabilizer tableau [96, 97] to the case of “quasi-stabilizers” (defined below), which permits efficient simulation of Lindblad dynamics for a certain class of non-Pauli stabilizers.
Finally, we comment on the values of correction rates required for these protocols and their feasibility on current erasure qubit platforms. We imagine a setup where each qubit gets acted on by $m$ gates per unit time for purposes of some desired quantum information processing task. We assume that each of these gates fails with probability $p_{g}$, resulting in a heralded error on the qubit. For small values of $p_{g}$, this can be modeled as an error process with rate $\eta_{g}=mp_{g}$. At the same time, the qubit is also affected by entangling gates that are applied during the measurement-feedback routine of our local correction protocol. The value of $\gamma$ shown in Fig. 1 indicates the rate at which the protocol must check for flags; however, the measurement and feedback gates are performed only when the classical erasure flags are in an appropriate correctable configuration. This rate is given by $\alpha\gamma$ ($\alpha<1$), with $\alpha\approx 0.03$ in the active phase of the $D_{4}$ model (see Fig. 17). If the measurement circuit fails with probability $p_{c}$, then the measurements during the local correction routine themselves contribute to the noise, with a rate $\eta_{c}=\alpha\gamma p_{c}$ [^5] Here, we are implicitly assuming that errors in the measurement-feedback routine have a qualitatively similar effect to applying a single-site Pauli error of the type analyzed in our model immediately following the measurement-feedback circuit.. Finally, qubits remaining in the idling state can decay from the computational subspace, leading to idling noise rate $\eta_{\text{idle}}$. Combining these contributions, we expect the total noise rate to be given by:
$\eta=\eta_{\text{idle}}+mp_{g}+\alpha\gamma p_{c}.$ (1)
As shown in Fig. 1(b), $D_{4}$ topological order can be stabilized provided that $\gamma/\eta=R\approx 200$. Since the rates must be non-negative, Eq. (1) can be satisfied only if $p_{c}<\frac{1}{R\alpha}$. For Yb-171 atoms [68], erasure qubits were recently demonstrated with $p_{g}\sim 0.02$ for two-qubit gates; taking $R\alpha\approx 10$ allows values of $p_{c}$ up to $0.1$, meaning that measurement-feedback must be achieved with less than 5 two-qubit gates in order to access the correction regime. In these systems $\eta_{\text{idle}}\sim 0.3s^{-1}$, while gate times are on the order of $\mu s$, suggesting that, provided the measurement error is sufficiently small, our protocol can easily control the idling errors.
@@ III Warm-up: Self-correcting 2d Toric code with heralded noise
We consider a honeycomb lattice defined on a two-dimensional torus, with a qubit (together with an $X$ and a $Z$ flag) placed on each edge. The periodic boundary conditions (PBC) on the $L_{x}\times L_{y}$ sized torus are implemented by making the following identifications for the position vector $\vec{r}$:
for $n_{x},n_{y}\in\mathbb{Z}$ and where $\vec{a},\vec{b}$, and $\vec{c}$ are the basis vectors of the honeycomb lattice, as shown in Fig. 2; these are related by $\vec{a}=\vec{b}+\vec{c}$. In the following, we set $L_{x}=L_{y}=L$, which results in a lattice with $L^{2}$ edges, $L^{2}/3$ plaquettes, and $2L^{2}/3$ vertices. Note that the Toric code can be defined on any 2D lattice; we choose the honeycomb lattice here to make contact with the $D_{4}$ model discussed later.
[…]
$B_{v}=\prod_{j\in^{*}v}Z_{j}\,,\qquad A_{p}=\prod_{k\in\partial p}X_{k}\,.$ (3)
@@ III.1 Noise model for unbiased heralded errors
We first introduce the noise model that we will use throughout this work, which consists of unbiased, heralded errors. To model a system with heralded errors, in addition to the qubits themselves, we also introduce a classical degree of freedom, which we refer to as an erasure “flag”. These flags are represented by a pair of occupation numbers $n^{f}_{X_{j}}$ and $n^{f}_{Z_{j}}$ which are associated with each qubit $j$ and can independently be in either occupied ($n^{f}_{X}=1,n^{f}_{Z}=1$) or unoccupied ($n^{f}_{X}=0,n^{f}_{Z}=0$) state. Here, we will set both $X$ and $Z$ flags on all qubits to be unoccupied $n^{f}_{X_{j}}=n^{f}_{Z_{j}}=0\,\forall\,j$ in the initial state.
@@
Here, $j$ labels the qubit index and $\mathsf{q_{x},q_{z}}\in\{0,1\}$. The operators $f^{+}_{X_{j}}(f^{-}_{X_{j}})$ raise $n^{f}_{X_{j}}$ from 0 to $1$ (lower $n^{f}_{X_{j}}$ from $1$ to 0). The relation $f^{+}_{X_{j}}f^{-}_{X_{j}}=n^{f}_{X_{j}}$ implies that the jump operator labeled by $\mathsf{q_{x}}=1(\mathsf{q_{x}}=0)$ acts on the site if an $X$ flag was already present (absent). The operators $f^{\pm}_{Z_{j}}$ act analogously on the $Z$ flags. Here